# Modeling Unknown Stochastic Dynamical System Subject to External Excitation

### v1

@@ 3.3.3 Normalizing Flow Model
$p_{\mathbf{Y}}(\mathbf{y})=p_{\mathbf{Z}}(\mathbf{f}(\mathbf{y}))|\det\mathbf{D}\mathbf{f}(\mathbf{y})|=p_{\mathbf{Z}}(\mathbf{f}(\mathbf{y}))|\det\mathbf{D}\mathbf{g}(f(\mathbf{y}))|^{-1},$
where $\mathbf{D}\mathbf{f}(\mathbf{y})=\partial\mathbf{f}/\partial\mathbf{y}$ is the Jacobian of $\mathbf{f}$ and $\mathbf{D}\mathbf{g}(\mathbf{z})=\partial\mathbf{g}/\partial\mathbf{z}$ is the Jacobian of $\mathbf{g}$. When the target complex distribution $p_{\mathbf{Y}}$ is given, usually as a set of samples of $\mathbf{Y}$, one chooses to find $\mathbf{g}$ from a parameterized family $\mathbf{g}_{\theta}$, where the parameter $\theta$ is optimized to match the target distribution. Also, to circumvent the difficulty of constructing a complicated nonlinear function $\mathbf{g}$, one utilizes a composition of (much) simpler diffeomorphisms: $\mathbf{g}=\mathbf{g}_{m}\circ\mathbf{g}_{m-1}\circ\cdots\circ\mathbf{g}_{1}$. It can be shown that $\mathbf{g}$ remains a diffeomorphism with its inverse $\mathbf{f}=\mathbf{f}_{1}\circ\cdots\circ\mathbf{f}_{m-1}\circ\mathbf{f}_{m}$. There exist a large amount of literature on normalizing flows. We refer interested reader to review articles [20, 32].
@@ 4 Numerical Examples
In this section, we present several numerical tests to demonstrate the performance of our proposed method. After presenting results for an Ornstein-Uhlenbeck (OU) process and a nonlinear SDE, we focus on nonlinear SDE systems for long-term predictions. These include stochastic a predator-prey model and a stochastic oscillator with double well potential. In both cases, we study very long-term predictions of the learned sFML models. In particular, for the stochastic oscillator, we utilize a periodic excitation signal that is known to generate the well-known “stochastic resonance” phenomenon.
In all the examples, the true SDE systems are known. However, the known SDEs are used only to generate the training data set (13). We solve the true systems by Euler-Maruyama method with a time step $\Delta=0.01$. The “initial conditions” $\mathbf{x}_{0}$ in (13) are sampled uniformly in a domain $I_{\mathbf{x}}$, specified in each example, and the excitations are local polynomials whose coefficients $\mathbf{\Gamma}_{0}$ are sampled in a domain specified for each example.
In our sFML model, Figure 1, the DNN $\mathbf{N}$ has 3 layers, each of which with 20 nodes, and utilizes $\tanh$ activation function. We employ cyclic learning rate with a base rate $3\times 10^{-4}$ and a maximum rate $5\times 10^{-4}$, $\gamma=0.99999$, and step size $10,000$. The cycle is set for every $40,000$ training epochs and with a decay scale $0.5$. A small weight decay of $0.01$ on the gradient updates is also used to help stablize the training. In our examples, the DNN training is usually conducted for $200,000\sim 300,000$ epochs.
@@ 4.1.1 OU with Drift Control
In Figure 2, we compare some sample trajectory pathes produced by the ground truth (left) and the learned sFML model (right), with an initial condition $x_{0}=2.0$ and a “new” control signal $\alpha(t)=\frac{1}{2}\sin(6t)$. We observe the two sets appear visually similar to each other. To further validate the sFML model prediction, we compute the mean and standard deviation of the solution averaged over $10,000$ trajectories. The sFML model predictions are shown in Figure 3, along with the reference ground truth. In Figure 4, we also show the comparison of the solution probability distributions at time $T=2,4,8$. We observe good agreement between the learned sFML model and the true model. This verifies that the sFML model indeed provides an accurate approximation in distribution.
@@ 4.1.2 Fully control
To examine the performance of the learned sFML model, we conduct a simulation with an initial condition $\mathbf{x}_{0}=1.0$ and excitations $\alpha(t)=\frac{1}{2}\sin(\frac{\pi}{2}t)$ and $\beta(t)=\frac{1}{10}e^{\cos(\pi t)}$. (Note that the excitations are not the Taylor polynomails in the training data set.) Some sample solution trajectories are shown in Figure 7. The mean and STD of the solution are shown in Figure 8. And in Figure 9, we also show the comparison of the probability distribution of the solution at $T=2,6,8$. We observe good agreement between the sFML model prediction and the gorund truth.
@@ 4.2 Nonlinear SDEs with Control
(23) $\left\{\begin{array}[]{l}\dot{x}_{t}=f(x_{t},y_{t},t)+\sigma_{1}\dot{W}_{1},\\ \dot{y}_{t}=-\mu(y_{t}-x_{t})+\sigma_{2}\dot{W}_{2},\end{array}\right.$
@@
Figure 12: Comparasion of probability distributions of Example 4.2 at $T=4,6,7,9$ with an initial condition $x_{0}=2.0$, $y_{0}=1.0$. To row: $x$; Bottom row: $y$.
@@ 4.3 Stochastic Predator-Prey Model
(24) $\left\{\begin{array}[]{l}\dot{x}_{t}=x_{t}-x_{t}y_{t}+u(t)+\sigma_{1}x_{t}\dot{W}_{1},\\ \dot{y}_{t}=-y_{t}+x_{t}y_{t}+\sigma_{2}y_{t}\dot{W}_{2},\end{array}\right.$
where ${W}_{1}$ and ${W}_{2}$ are independent Brownian motions, and $\sigma_{1}=\sigma_{2}=0.05$. The training data are generated by simulating $120,000$ solutoin samples for one step $\Delta=0.01$, from initial conditions in $I_{\mathbf{x}}=[0.1,0.35]\times[0.2,5.5]$ and under exicitations of 2nd-degree Taylor polynomials whose coefficients are from $[0.01,4.2]\times[-1.5,1.5]\times[-0.7,0.7]$.
@@ 4.4 Stochastic Resonance
Finally, we consider the following SDE with a double-well potential and excitation,
@@
where $\sigma=0.25$ is a parameter, and $u(t)$ is the excitation. When $V=0$, there is no excitation to the system. The solution would exhibit random transition between two metastable states $x=-1$ and $x=1$. The transition probability depends on the parameters $\sigma$. When $V\neq 0$, an excitation is exerted to the system. If the excitation is periodic, under the right circumstance the random transtion between the two metastable states becomes synchorized with the perodicity of the exication, resulting in the so-called stochastic resonance, cf., [4, 2, 3].
@@
Once the sFML model is trained, we conduct system prediction under various excitations. In particular, we choose $u(t)=V\cos(\omega t)$, with $V=0.12$ and $\omega=0.001$. These parameters are chosen according to [4], to ensure the occurrence of stochastic resonance. An exceptionially long-term system prediction is conducted by the sFML model, for time up to $T=40,000$. The result is shown in the top of Figure 15, where we also plotted the (rescaled) periodic excitation in light grey line in the background. We can clearly observe the synchonization between the random transition and the periodic excitation — the stochastic resonance. For reference, we also conduct the sFML system prediction with $V=0$, i.e., no excitation. The solution, shown in the bottom of Figure 15, exhibits the expected random transition between the two metastable states. We shall emphsize that in this case the transition probability is very small, $O(10^{-5})$. The learned sFML model is capable of capturing such a small probability event. We shall remark again that the training data are pairwise data separated by one time step. Thus, none of the (long-term) system behaviors can be observed in the training data.
@@ 5 Conclusion
In this paper, we presented a general numerical framework for modeling unknown nonautonomous stochastic systems by using observed trajectory data. To overcome the difficulties brought by the external time-dependent inputs, we transfer the original system into a local parametric stochastic system. We accomplished this by locally parameterizing the time-dependent external inputs on several discrete time points. The resulting stochastic system is then driven by a stationary parametric stochastic flow map. A normalizing flow model is devised to approximate the parametric stochastic flow map. By using a comprehensive set of numerical examples, we demonstrated that the proposed approach is effective and accurate in modeling a variety of unknown stochastic systems. The learned model can conduct expectational long-term system, subject to arbitrary external excitations that are not contained in the training data.

### v2

@@ 3.3.3 Normalizing Flow Model
$p_{\mathbf{Y}}(\mathbf{y})=p_{\mathbf{Z}}(\mathbf{f}(\mathbf{y}))|\det\mathbf{D}\mathbf{f}(\mathbf{y})|=p_{\mathbf{Z}}(\mathbf{f}(\mathbf{y}))|\det\mathbf{D}\mathbf{g}({\color[rgb]{0,0,0}\mathbf{f}}(\mathbf{y}))|^{-1},$
where $\mathbf{D}\mathbf{f}(\mathbf{y})=\partial\mathbf{f}/\partial\mathbf{y}$ is the Jacobian of $\mathbf{f}$ and $\mathbf{D}\mathbf{g}(\mathbf{z})=\partial\mathbf{g}/\partial\mathbf{z}$ is the Jacobian of $\mathbf{g}$. When the target complex distribution $p_{\mathbf{Y}}$ is given, usually as a set of samples of $\mathbf{Y}$, one chooses $\mathbf{g}$ from a parameterized family $\mathbf{g}_{\theta}$, where the parameter $\theta$ is optimized to match the target distribution. Also, to circumvent the difficulty of constructing a complicated nonlinear function $\mathbf{g}$, one utilizes a composition of (much) simpler diffeomorphisms: $\mathbf{g}=\mathbf{g}_{m}\circ\mathbf{g}_{m-1}\circ\cdots\circ\mathbf{g}_{1}$. It can be shown that $\mathbf{g}$ remains a diffeomorphism with its inverse $\mathbf{f}=\mathbf{f}_{1}\circ\cdots\circ\mathbf{f}_{m-1}\circ\mathbf{f}_{m}$. There exist a large amount of literature on normalizing flows. We refer interested reader to review articles [25, 41].
@@ 4 Numerical Examples
In this section, we present several numerical tests to demonstrate the performance of our proposed method. After presenting results for learning an Ornstein-Uhlenbeck (OU) process and a nonlinear SDE, we focus on nonlinear SDE systems for long-term predictions. These include stochastic a predator-prey model and a stochastic oscillator with double well potential. In both cases, we study very long-term predictions of the learned sFML models. In particular, for the stochastic oscillator, we utilize a periodic excitation signal that is known to generate the well-known “stochastic resonance” phenomenon. At the end of this section, we present learning of a stochastic reaction network and a stochastic partial diferential equation (SPDE). The stochastic reaction network is a non-Gaussian stochastic jump process driven by its inherent stochasticity that is not Brownian motion. It demonstrates the applicability of the proposed method to general stochastic system (1) beyond the classical SDE (3). The SPDE example demonstrates the applicability of the method to a modestly high dimension $d=30$.
In all the examples, the true stochastic systems are known. However, the true stochastic systems are used only to generate the training data set (15). When the true systems follow the SDE form (3), we solve them by Euler-Maruyama method with a time step $\Delta=0.01$. The “initial conditions” $\mathbf{x}_{0}$ in (15) are sampled uniformly in a domain $I_{\mathbf{x}}$, specified in each example, and the excitations are local polynomials whose coefficients $\mathbf{\Gamma}_{0}$ are sampled in a domain specified for each example.
In our sFML model, Figure 1, the DNN $\mathbf{N}$ has 3 layers, each of which with 20 nodes, and utilizes $\tanh$ activation function. We employ cyclic learning rate with a base rate $3\times 10^{-4}$ and a maximum rate $5\times 10^{-4}$, $\gamma=0.99999$, and step size $10,000$. The cycle is set for every $40,000$ training epochs and with a decay scale $0.5$. A small weight decay of $0.01$ on the gradient updates is also used to help stabilize the training. In our examples, the DNN training is usually conducted for $200,000\sim 300,000$ epochs.
For validation tests, we conduct system predictions by the learned FML for time domains well beyond that of the training data set. We employ mostly periodic excitation signals to ensure that the long-term system behavior remains “interesting”, that is, it does not either decay to zero or blow up. We remark that periodicity of the exciation is not required in the method, as it was never assumed in the derivation of the method.
@@ 4.1.1 OU with Drift Control
In Figure 2, we compare some sample trajectory paths produced by the ground truth (left) and the learned sFML model (right), with an initial condition $x_{0}=2.0$ and a “new” control signal $\alpha(t)=\frac{1}{2}\sin(6t)$. We observe the two sets appear visually similar to each other. To further validate the sFML model prediction, we compute the mean and standard deviation of the solution averaged over $10,000$ trajectories. The sFML model predictions are shown in Figure 3, along with the reference ground truth. In this figure, lines represent the mean of sample trajectories, while shadow bands depict one standard deviation above and below the mean curve. In Figure 4, we also show the comparison of the solution probability distributions at time $T=2,4,8$. We observe good agreement between the learned sFML model and the true model. This verifies that the sFML model indeed provides an accurate approximation in distribution.
@@ 4.1.2 Fully control
To examine the performance of the learned sFML model, we conduct a simulation with an initial condition $\mathbf{x}_{0}=1.0$ and excitations $\alpha(t)=\frac{1}{2}\sin(\frac{\pi}{2}t)$ and $\beta(t)=\frac{1}{10}e^{\cos(\pi t)}$. (Note that the excitations are not the Taylor polynomials in the training data set.) Some sample solution trajectories are shown in Figure 7. The mean and STD of the solution are shown in Figure 8. And in Figure 9, we also show the comparison of the probability distribution of the solution at $T=2,6,8$. We observe good agreement between the sFML model prediction and the ground truth.
@@ 4.2 Nonlinear SDEs with Control
(25) $\left\{\begin{array}[]{l}dx_{t}=f(x_{t},y_{t},t)dt+\sigma_{1}dW_{1},\\ dy_{t}=-\mu(y_{t}-x_{t})dt+\sigma_{2}dW_{2},\end{array}\right.$
@@
Figure 12: Comparasion of probability distributions of Example 4.2 at $T=4,6,7,9$ with an initial condition $x_{0}=2.0$, $y_{0}=1.0$. Top row: $x$; Bottom row: $y$.
@@ 4.3 Stochastic Predator-Prey Model
(26) $\left\{\begin{array}[]{l}dx_{t}=\left[x_{t}-x_{t}y_{t}+u(t)\right]dt+\sigma_{1}x_{t}dW_{1},\\ dy_{t}=(-y_{t}+x_{t}y_{t})dt+\sigma_{2}y_{t}dW_{2},\end{array}\right.$
where ${W}_{1}$ and ${W}_{2}$ are independent Brownian motions, and $\sigma_{1}=\sigma_{2}=0.05$. The training data are generated by simulating $120,000$ solution samples for one step $\Delta=0.01$, from initial conditions in $I_{\mathbf{x}}=[0.1,0.35]\times[0.2,5.5]$ and under excitations of 2nd-degree Taylor polynomials whose coefficients are from $[0.01,4.2]\times[-1.5,1.5]\times[-0.7,0.7]$.
@@ 4.4 Stochastic Resonance
In this section, we consider the following SDE with a double-well potential and excitation,
@@
where $\sigma=0.25$ is a parameter, and $u(t)$ is the excitation. When $V=0$, there is no excitation to the system. The solution would exhibit random transition between two metastable states $x=-1$ and $x=1$. The transition probability depends on the parameters $\sigma$. When $V\neq 0$, an excitation is exerted to the system. If the excitation is periodic, under the right circumstance the random transition between the two metastable states becomes synchronized with the periodicity of the excitation, resulting in the so-called stochastic resonance, cf., [5, 3, 4].
@@
Once the sFML model is trained, we conduct system prediction under various excitations. In particular, we choose $u(t)=V\cos(\omega t)$, with $V=0.12$ and $\omega=0.001$. These parameters are chosen according to [5], to ensure the occurrence of stochastic resonance. An exceptionally long-term system prediction is conducted by the sFML model, for time up to $T=40,000$. The result is shown in the top of Figure 15, where we also plotted the (rescaled) periodic excitation in light grey line in the background. We can clearly observe the synchronization between the random transition and the periodic excitation — the stochastic resonance. For reference, we also conduct the sFML system prediction with $V=0$, i.e., no excitation. The solution, shown in the bottom of Figure 15, exhibits the expected random transition between the two metastable states. We shall emphasize that in this case the transition probability is very small, $O(10^{-6})$. The learned sFML model is capable of capturing such a small probability event. We shall remark again that the training data are pairwise data separated by one time step. Thus, none of the (long-term) system behaviors can be observed in the training data. In Figure 16, we also show the comparison of reference and learned density functions of the test trajectory at time $T=10,000,20,000,30,000,40,000$. We observe that the sFML model exhibits good accuracy in these predictions.
@@
Figure 16: Comparasion of probability distributions of Example 4.4 at $T=10,000,20,000,30,000,40,000$ with the periodic excitation $u(t)=V\cos(\omega t)$, $V=0.12$ and $\omega=0.001$.
4.5 A Gene Expression Model
We consider a gene expression model in the form of biochemical reaction network [53, 6], described as Example 2 in Section 2,
(28) $\begin{split}\emptyset&\xrightarrow{k(t)}M\\ M&\xrightarrow{k_{\text{s}}}M+P\\ M&\xrightarrow{k_{\text{dm}}}\emptyset\\ P&\xrightarrow{k_{\text{dp}}}\emptyset,\end{split}$
where there are 2 species: $M$ represents mRNA and $P$ represents protein. In this two-stage reaction network model, mRNAs are transcribed in a time-dependent rate $k(t)$ and proteins are translated at a rate of $k_{\text{s}}$ off each mRNA molecule. Meanwhile, mRNAs and proteins degrade at rate $k_{\text{dm}}$ and $k_{\text{dp}}$. The dynamical behavior of state variables $(M,P)$ is usually modeled as a stochastic jump process due to the stochasticity caused by low number of molecules. Stochastic Simulation Algorithm (SSA, c.f. [20]) is an exact Monte Carlo method designed for producing realizations of such processes. Though the transitional probability of these processes is non-Gaussian and nearly impossible to be explicitly written, we demonstrate in this example that the proposed sFML method is capable of learning and predicting such systems.
We simulate the system using Modified Next Reaction method [1], which is a variant of the original SSA. The training data are generated via simulating the system up to a short time $\Delta=0.1$ with randomly sampled initial conditions for $M\leq 10$ and $P\leq 400$. Taylor polynomials of degree $2$ are used to conduct local approximation of the excitation signal $k(t)$ within the $\Delta$ time interval. The 3 parameters for $\mathbf{\Gamma}_{0}$ are sampled uniformly from $[0,40]\times[-5.3,5.3]\times[-0.7,0.7]$. For other parameters, we fix $\Delta=0.1$, $k_{\text{s}}=500$, $k_{\text{dm}}=20$, $k_{\text{dp}}=5$. (Note that these parameters are only used to generate the training data. They are not included in the FML model.)
For the learned sFML model, we conduct a system prediction with an initial condition $(M_{0},P_{0})=(1,1)$ and $k(t)=20+20\sin(\frac{\pi x}{12})$ up to time $T=240$. To test the learning accuracy, we fix $(\mathbf{x}_{0},\mathbf{\Gamma}_{0})$, generate $10,000$ samples of $\mathbf{z}$ for $\mathbf{G}_{\Delta}(\mathbf{x}_{0},\mathbf{z};\mathbf{\Gamma}_{0})$ to compare the conditional distribution. In Figure 17, we show the 2D histogram for 2 scenarios: $\mathbf{x}_{0}=(2,133)$, $\mathbf{\Gamma}_{0}=(30,-4.535,-0.335)$ and $\mathbf{x}_{0}=(5,287)$, $\mathbf{\Gamma}_{0}=(39.319,1.355,-0.664)$. The sFML model achieves high accuracy in this non-Gaussian transitional probability distribution case.
Figure 17: 2D histogram of (non-Gaussian) conditional distribution $\mathbf{G}_{\Delta}(\mathbf{x}_{0},\mathbf{z};\mathbf{\Gamma}_{0})$ with fixed $(\mathbf{x}_{0},\mathbf{\Gamma}_{0})$ of Example 4.5. Left 2 figures: reference and learned distribution for $\mathbf{x}_{0}=(2,133)$ and $\mathbf{\Gamma}_{0}=(30,-4.535,-0.335)$; Right 2 figures: reference and learned distribution for $\mathbf{x}_{0}=(5,287)$ and $\mathbf{\Gamma}_{0}=(39.319,1.355,-0.664)$.
Figure 18: Sample trajectory for Example 4.5 with initial condition $(M_{0},P_{0})=(1,1)$ and $k(t)=20+20\sin(\frac{\pi x}{12})$. Left: a reference sample trajectory; Right: a learned sample trajectory.
Figure 19: Mean and standard deviation (STD) for Example 4.5 with initial condition $(M_{0},P_{0})=(1,1)$ and $k(t)=20+20\sin(\frac{\pi x}{12})$. Left: $M$; Right: $P$.
Figure 20: Comparasion of probability distributions of Example 4.5 at $T=50,60,78,110$ with initial condition $(M_{0},P_{0})=(1,1)$ and $k(t)=20+20\sin(\frac{\pi x}{12})$. Top row: $M$; Bottom row: $P$.
In Figure 18, we show one trajectory produced by SSA (left) and the learned sFML model (right). These trajectories reveal similar patterns. To further validate the long time accuracy, we compute the mean and standard deviation with $10,000$ trajectories and show the results in Figure 19. Moreover, the probability distributions at time $T=50$,$60$,$78$, $110$ are shown in Figure 20. We observe good agreement between the sFML model predictions and the ground truth.
4.6 A Stochastic Heat Equation with Source
In this section, we consider learning a one-dimensional stochastic heat equation driven by space-time noise
(29) $\partial_{t}u=\varepsilon\Delta u+f(x,t)+\sigma\xi(x,t),\qquad(x,t)\in(0,2\pi)\times(0,T],$
where $\xi(x,t)$ is a space-time white noise, $f(x,t)=\alpha(t)e^{-\frac{(x-p)^{2}}{q^{2}}}$, $\varepsilon$, $p$, $q$ and $\sigma$ are model parameters, $\alpha(t)$ is the excitation. Periodic boundary condition is prescribed and $\varepsilon=0.1$.
We conduct the sFML modeling in modal space, i.e., in Fourier space. Assume the trajectory data of (29) are observed at grid points $x_{j}=jh$, $j=0,1,...,N-1$ where $h=2\pi/N$. These observational data are then projected onto a finite dimensional space $\mathbb{V}_{N}=\text{span}\{e_{k}\}_{k=1}^{N}$, where $\{e_{k}\}_{k=1}^{\infty}$ is an orthogonal basis of $\mathbb{V}=L^{2}_{\text{per}}(0,2\pi)$. Thus, the solution $u(x,t)$ can be approximated by
(30) $u_{N}(x,t)=\sum_{i=0}^{N}c_{k}(t)e_{k}(x),$
where we use the Fourier basis for $e_{k}$’s and $c_{k}$’s are the Fourier coefficients. We solve (29) with Fourier collocation method with $N=30$. We then apply the proposed sFML method to learn the dynamical behavior of the coefficients $\mathbf{c}=[c_{0},\dots,c_{N}]^{T}$, where the excitation signal $\alpha(t)$ is represented as 2nd degree Taylor polynomial with the coefficient $\mathbf{\Gamma}_{0}\in(-1.2,1.2)\times(-3.5,3.5)\times(-5,5)$. A total of $120,000$ trajectory pairs are used in the training data set (15), where the time step $\Delta=0.05$.
For validation, we conduct the sFML simulation for $p=1.0$, $q=1.0$, $\sigma=0.05$, and with $\alpha(t)=\sin t$. The initial condition is set as $u(x,0)=\exp(-\sin^{2}x)-1$. In the top of Figure 21, we compute the temporal evolution of the mean and standard deviation of the solution as several locations, averaged over $10,000$ trajectories. In the bottom of Figure 21, we present the solution at several different time instances $T=0.5,2.5,5,8$. In Figure 22, we also show the comparison of the probability distribution of the solution at a few arbitrarily chosen spatial-temporal locations, $(x,t)=(0.7,2),(1.3,4),(3.8,6),(5.5,8)$. We observe good agreement between the sFML model prediction and the ground truth.
Figure 21: Mean and standard deviation (STD) of Example 29 with initial condition $u(x,0)=\exp(-\sin^{2}x)-1$ and $\alpha(t)=\sin t$. Upper: Comparison for $u(x,\cdot)$ with $x=0,\pi/2,\pi,3\pi/2$. Lower: Comparison for $u(\cdot,T)$ with $T=0.5,2.5,5,8$.
Figure 22: Comparasion of probability distributions of Example 29 at $(x,t)=(0.7,2),(1.3,4),(3.8,6),(5.5,8)$ with initial condition $u(x,0)=\exp(-\sin^{2}x)-1$ and $\alpha(t)=\sin t$.
@@ 5 Conclusion
In this paper, we presented a general numerical framework for modeling unknown nonautonomous stochastic systems by using observed trajectory data. To overcome the difficulties brought by the external time-dependent inputs, we transfer the original system into a local parametric stochastic system. We accomplished this by locally parameterizing the time-dependent external inputs on several discrete time points. The resulting stochastic system is then driven by a stationary parametric stochastic flow map. A normalizing flow model is emplooyed to approximate the parametric stochastic flow map. By using a comprehensive set of numerical examples, we demonstrated that the proposed approach is effective and accurate in modeling a variety of unknown stochastic systems. One of the key features of the method is that it is able to produce learned models that can conduct accurate system predictions under external excitations not in the training data and for exceptionally long-term well beyond the time domain of the training data. Although the method works remarkably well for the variety of examples in the paper, we caution that its rigorous numerical analysis such as error estimate is still lacking, largely due to the lack of fudamental results on DNNs. Also, the performance for high dimensional stochastic systems is unclear. These issues are being pursued at this moment and the results will be reported in future studies.